OpenStudy (anonymous):

Use a transformation to solve the equation. s – 14.2 = –18.7 A. s = –32.9 B. s = –4.5 C. s = 4.5 D. s = 32.9Use a transformation to solve the equation. s – 14.2 = –18.7 A. s = –32.9 B. s = –4.5 C. s = 4.5 D. s = 32.9

OpenStudy (anonymous):

s – 14.2 = –18.7 s = -18.7 + 14.2 s= -4.5

OpenStudy (studytheworld):

s = -4.5.

OpenStudy (kaylardz0405):

#2 18.7+14.2

OpenStudy (anonymous):

NO DIRECT ANSWERS @StudyTheWorld!!!

OpenStudy (anonymous):

Do u guys know who @Anonymous_Hacker is

OpenStudy (kaylardz0405):

no

OpenStudy (anonymous):

@KaylaRdz0405 It is actually -18.7 + 14.2

OpenStudy (anonymous):

This is irrelevant.

OpenStudy (kaylardz0405):

oh sorry

OpenStudy (anonymous):

Its fine. :) Medal anyway. :)